QUESTION: Find the oxidation number of carbon in CH3
OH
.

Grade:

4 Answers

Vikas TU
3 years ago
Oxigation no of O and H are -2 and +1 
x+3(1)+ (-2)+1 = 0 
x+3-2+1 = 0 
x = +2 
Oxidation no is +2 
good luck
ankit singh
3 years ago
X+ 3- 2+ 1=0
Solve for x, we get x=2 which is oxidation number and valency is 4.. thsta the correct answer
ajay
3 years ago
Oxigation no of O and H are -2 and +1 
x+3(1)+ (-2)+1 = 0 
x+3-2+1 = 0 
x = +2 
Oxidation no is +2 Oxigation no of O and H are -2 and +1 
x+3(1)+ (-2)+1 = 0 
x+3-2+1 = 0 
x = +2 
Oxidation no is +2 
X+ 3- 2+ 1=0
Solve for x, we get x=2 which is oxidation number and valency is 4.. thsta the correct answer
 
ajay
3 years ago
Oxigation no of O and H are -2 and +1 
x+3(1)+ (-2)+1 = 0 
x+3-2+1 = 0 
x = +2 
Oxidation no is +2 
Oxigation no of O and H are -2 and +1 
x+3(1)+ (-2)+1 = 0 
x+3-2+1 = 0 
x = +2 
Oxidation no is +2